OpenStudy (anonymous):

Is the solution to the radical function sqrtx^2-3x=2 x=1, x=4 or both???? I'll give medals! Please help.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (irishboy123):

\(\large \sqrt{x^2-3x}=2\)??? if so, \(\large x^2-3x=4, \, (x+1)(x-4) = 0\)